# The Evolution and Future Perspectives of Artificial Intelligence Generated Content

### v1

@@ Abstract
Artificial intelligence generated content (AIGC), a rapidly advancing technology, is transforming content creation across domains, such as text, images, audio, and video. Its growing potential has attracted more and more researchers and investors to explore and expand its possibilities. This review traces AIGC’s evolution through four developmental milestones-ranging from early rule-based systems to modern transfer learning models-within a unified framework that highlights how each milestone contributes uniquely to content generation. In particular, the paper employs a common example across all milestones to illustrate the capabilities and limitations of methods within each phase, providing a consistent evaluation of AIGC methodologies and their development. Furthermore, this paper addresses critical challenges associated with AIGC and proposes actionable strategies to mitigate them. This study aims to guide researchers and practitioners in selecting and optimizing AIGC models to enhance the quality and efficiency of content creation across diverse domains.
@@ I Introduction
As shown in Figure 1, many models and algorithms have been developed along the progressive trajectory of AIGC, with each technological advancement built upon the previous ones. Fresh endeavors persist in unfolding, leading to great tools [7][1] and systems [8][9] that support sustainable and adaptable content production across various sectors. Part of the reason for so much interest in this development is the complexity of the problem and its many associated challenges. While data have been a fundamental drive to advance digital transformation across industries, its collection is often insufficient [10][11] or sometimes impossible [12]. Data annotation is rather time-consuming and costly, leading to the available public datasets often being on a small scale. Examples include breast pathological images [13], the MNIST dataset [14] for handwritten digits, and the CoNLL-2003 dataset [15] for named entity recognition. When significant manpower is added, it becomes harder to maintain consistency in tone, style, and quality. These inherent problems highlight the need for rapid, scalable, and consistent content production methods.
@@
Content generation approaches vary in terms of methods, and more importantly, in terms of limitations. Some have experimented with simple text generation to simulate human-like language interactions via rule-based methods [16][17], while others capture statistical dependency between text elements in a sequence to predict the next element [18]. These earlier attempts rely heavily on either experts’ rules established prior or a set of probabilistic measures, scoping their capabilities within predefined scenarios. With the involvement of deep learning, the performance of generative models has been substantially improved, unlocking opportunities for a vast range of applications [4]. Meanwhile, the critical need for trustworthiness in these models is also raised [19].
@@
The contributions of this paper are twofold. First, this paper offers an in-depth review of the historical development of AIGC, mapping its journey from early rule-based systems to advanced neural network methodologies. More importantly, this review addresses both technological advancements and challenges such as scalability and ethical dilemmas.
Second, unlike most of existing review papers that focus on either a specific application domain, such as text-to-3D generation (Li et al. 2023) [20][21], or a particular type of methodology such as Video GANs (Aldausari et al. 2023) [22][23], our paper our work resembles [24] in its exploration of the historical evolution of AIGC but with even broader coverage. More importantly, this paper integrates a wider range of methodologies and applications into a cohesive framework that spans multiple phases of AIGC development. A unique contribution of this work is the use of a consistent example throughout, which not only enables readers to follow the evolution of AIGC methods but also makes it easier to understand and compare the differences among various approaches within a practical context. While we share the same view of the challenges facing AIGC, such as bias, ethical concerns, and scalability of models [25][26], this paper extends the discussions by proposing research directions to tackle these issues. With these directions outlined, this paper contributes to the ongoing discourse on responsible AIGC development in real-world applications.
The rest of the paper is organized as follows. Section II provides the criteria for literature inclusion in this review. A detailed exploration of AIGC is given in Section III, presenting its core characteristics and advantages such as automation, consistency, and efficiency. Section IV delves into the early rule-based systems that lay the foundation for AIGC, followed by a discussion of the evolution towards statistical methods, deep learning models, and the integration of transfer and pre-trained models in Sections V, VI, and VII respectively. Each section examines technological advancements and applications of these methodologies in real-world scenarios. Section VIII addresses the potential challenges and limitations of current AIGC technologies. Finally, Section IX concludes the review by summarizing the comprehensive insights gathered throughout the paper and suggesting future research directions.
@@ II Criteria for Inclusion in the Literature Review
To ensure comprehensive coverage of recent and seminal research in the field of AIGC, the literature search process begins with several electronic databases, including IEEE Xplore, Scopus, ArXiv, and Google Scholar with the keywords “AIGC,” “artificial intelligence generated content,” “machine learning content generation,” and “automated content creation”. While a time constraint is not imposed on published research, we make sure to incorporate the latest advancements in AIGC technology and applications. Non-English papers, short communications, and non-peer-reviewed articles are excluded to ensure the scientific rigor of the review. Each selected paper has to make a demonstrable contribution to the development, application, or evaluation of AIGC systems. The initial search yields more than 300 potential articles, and after screening for relevance to core themes of AIGC, the pool of relevant papers is narrowed to 200. The outcome of this meticulous search process is summarized in Figure 2, which displays the distribution of articles across journals over the years, illustrating the focus areas and publication trends in AIGC research.
@@ III General AIGC Technique Framework
AIGC is an innovative way of autonomously generating diverse forms of content by algorithms, machine learning models, or other artificial intelligence techniques. The main characteristic of AIGC is its reliance on data input and pre-generated models to produce content, requiring less or no human intervention. Compared to manual creation, AIGC has distinguished advantages in terms of automation, scale, and efficiency [27]. It also ensures consistency, customization capability [28], and cost-effectiveness [29]. These strengths have continuously motivated researchers [30] to advance AIGC technologies for decades, leading to a variety of applications as shown in Table I. Beginning with basic text generation [31], AIGC has now expanded to include video [6], audio [32], and 3D model [33] generation for applications in fields such as education [34], healthcare [35], entertainment [36], and many others.
@@
The flow of an AIGC system begins with user input that goes through a pre-processing phase to covert from its raw, naturally human-understandable form into a structured, machine-interpretable format [68]. This usually involves techniques such as natural language processing (NLP) [69] for text data or feature extraction for visual [70] or auditory data, with methods like decision trees and pattern matching bridging input rules to algorithms.
@@
These sequential steps, each serving a unique role, contribute collaboratively to the robustness and efficacy of the content generation process. A deep understanding of these steps in each developmental milestone of AIGC helps us gain insights into where AIGC has been and where it is heading. To that end, this paper uses the same example of “generate a research question with the key words: artificial intelligence, healthcare, and ethical implications” across all milestones. This unified reference point not only demonstrates the unique advantages of the methodologies at each milestone, such as the transparency of rule-based systems, the data-driven nature of statistical methods, the automation capabilities of deep learning, and the flexibility of transfer learning but also makes comparisons between approaches more straightforward and clear. Ultimately, the paper aims to assist researchers in better grasping each method’s operational process, enabling them to choose a proper one for their practical applications or combine different ones for optimal results.
@@ IV Early Rule-based Systems
The roots of AIGC can be traced back to the 1950s [73][74] when early rule-based systems lays the foundation for its development. These systems employ a set of expert rules defined a priori to guide content generation. Figure 4 presents the specific framework for rule-based systems within the context of AIGC discussed in the previous section.
@@ IV-A Fundamentals
Early rule-based systems primarily focus on texts, using NLP techniques to enable computers to understand human inputs, in which the preprocessing step is attributed to ‘parsing’. The common methods include tokenization [69], stemming, stop word removal, rule-based parsing, and entity recognition [17]. Tokenization breaks down the input data into small pieces, like words. Doing so converts a continuous stream of text into a manageable sequence of tokens, making it easier for algorithms to interpret. Stemming reduces words to their basic form, like ’running’ to ’run’. This ensures consistency across similar words, reducing complexity with fewer unique words. In the step of stop work removal, common yet nonessential words like ’the’ and ’is’ are removed. This step helps the system focus on meaningful words only and reduces noise in the data. Rule-based parsing applies grammar rules to understand sentence structure, aiding in the interpretation of word relationships and the comprehension of text syntax and semantics. In addition, Entity recognition helps spot and categorize keywords, enabling the quick identification of important information such as the text’s main subjects. These steps streamline the complexity of natural language, making it more structured and manageable for analysis and encoding.
@@
The role of post-processing in the rule-based phase is pretty straightforward. It usually consists of some content enrichment processes, such as sentence re-organization, and spelling and grammar corrections, to improve the clarity and fluency of generated results. The decision to enhance the consistency or diversity and depth of the language used is made through content formatting and lexical enrichment.
@@ IV-B Example
Early rule-based systems find applications in text generation. For example, Eliza [16] is designed as a chatbot for psychotherapist conversations, SHRDLU [17] advances semantic analysis, and Tale-Spin [78] can create poetry and stories. Here, we detail how Eliza responds to the user input “generate a research question with the key words: artificial intelligence, healthcare, and ethical implications” to showcase the operations of early rule-based systems.
@@ V-A Fundamentals
Gaussian Mixture Models (GMM) 1980s L.P. Hansen[82]
@@
In the realm of statistical language processing, N-gram models serve as a straightforward yet effective method for predicting the next item in a sequence based on the preceding N items. They are particularly valuable for tasks such as lexical selection and modifier ordering [85] due to their ability to capture local context. Yet N-gram is effective for text generation and speech recognition [86], it often faces several limitations, especially when dealing with long-range dependencies and language ambiguity, owing to their fixed context window. To address these limitations, researchers develop Hidden Markov Models (HMMs), which build upon N-grams by incorporating hidden states, allowing for more flexible sequence modeling. HMMs are crucial for handling temporal variations, making them indispensable in tasks like part-of-speech tagging and speech recognition[87] [88][89]. Nonetheless, HMMs are still constrained by their assumptions about state independence and often require extensive training data to achieve high accuracy.
@@ VI Deep Learning Methods
Neural networks and deep learning models mark a major leap forward in AIGC [94], overcoming the limitations of statistical methods by better managing the growing volume, dimensionality, and complexity of modern data, especially unstructured data such as images, text, and speech. Mimicking the structure of the human brain, neural networks consist of layers of interconnected nodes (neurons). As data flows through the layers, the networks conduct a step-by-step inference process, automatically recognizing patterns and making predictions from input to output. Extended from this foundational structure, deep learning uses convolutional neural networks (CNNs), in which the nodes at each layer are organized into overlapping clusters so that each cluster feeds data to multiple nodes in subsequent layers. This design enhances the network’s capability to process high-dimensional data to learn and capture complex relationships more effectively. As shown in Figure 9, deep-learning-based AIGC shares similarities with statistical methods in terms of feature engineering for user input. What sets them apart is the process of selecting, modifying, and creating features. Deep learning automates that process through training on diverse and large datasets while statistical methods require human intervention. Notably, the quality and diversity of datasets are crucial for training deep learning models, as the data needs to represent a wide range of scenarios within the problem’s scope. Thus, data augmentation is often applied to enhance both the size and variability of the training dataset, improving the model’s generalization ability. In deep learning, the backbone architectures include Recurrent Neural Networks (RNNs), CNNs, and Transformer [95] as detailed in the following subsections.
@@ VI-C Transformer
As shown in Figure 10, the transformer consists of two main parts: multiple layers $j$ (i.e., $j=N$) of an encoder and a decoder that have very similar structures but serve different purposes. The former is responsible for encoding the input sequence into a context-rich representation; while the latter is used to generate new information in a sequence based on the given input and its encoded representation.
Before the input sequence is fed into the encoder, it is first tokenized, similar to that in early rule-based and statistical methods. What sets the transformer apart is the embedding process, mapping each token into a dense vector representation with a dimension of $d_{\text{model}}$. Note that such dense vectors are generated from an embedding matrix, either derived from a pre-trained model (e.g., BERT[98], GPT[7]) or trained from scratch during the model’s training stage. As stated earlier, the transformer model processes the entire input in parallel, it does not have an inherent understanding of the sequence order of the tokens. Therefore, positional encoding is essential, as it is ultimately added to the input embedding to help the model gain the necessary sequence order information. While sine and cosine are the most traditional and widely used forms of positional encoding, many other approaches are available, such as learned embeddings[98] and rotary positional embeddings[99], and their choice highly depends on the model architecture and the task requirements.
@@
While initially introduced for NLP tasks, the modularity and extensibility of the transformer model have made it a popular and powerful tool across a wide range of applications. Beyond excelling in text generation, it has recently been successfully extended to the area of computer vision, such as Vision Transformer [100], showcasing its effectiveness in vision generation. Many large language models, such as BERT [98] and GPT [101][102][7], are built upon its core architecture.

### v2

@@ Abstract
Artificial intelligence generated content (AIGC), a rapidly advancing technology, is transforming content creation across domains, such as text, images, audio, and video. Its growing potential has attracted more and more researchers and investors to explore and expand its possibilities. This review traces AIGC’s evolution through four developmental milestones, ranging from early rule-based systems to modern transfer learning models, within a unified framework that highlights how each milestone contributes uniquely to content generation. In particular, the paper employs a common example across all milestones to illustrate the capabilities and limitations of methods within each phase, providing a consistent evaluation of AIGC methodologies and their development. Furthermore, this paper addresses critical challenges associated with AIGC and proposes actionable strategies to mitigate them. This study aims to guide researchers and practitioners in selecting and optimizing AIGC models to enhance the quality and efficiency of content creation across diverse domains.
@@ I Introduction
As shown in Figure 1, many models and algorithms have been developed along the progressive trajectory of AIGC, with each technological advancement built upon the previous ones. Fresh endeavors persist in unfolding, leading to great tools [1] and systems [7][8] that support sustainable and adaptable content production across various sectors. Part of the reason for so much interest in this development is the complexity of the problem and its many associated challenges. While data has been a fundamental drive to advance digital transformation across industries, its collection is often insufficient [9][10] or sometimes impossible [11]. Data annotation is rather time-consuming and costly, leading to the available public datasets often being on a small scale. Examples include breast pathological images [12], the MNIST dataset [13] for handwritten digits, and the CoNLL-2003 dataset [14] for named entity recognition. When significant manpower is added, it becomes harder to maintain consistency in tone, style, and quality. These inherent problems highlight the need for rapid, scalable, and consistent content production methods.
@@
Content generation approaches vary in terms of methods, and more importantly, in terms of limitations. Some have experimented with simple text generation to simulate human-like language interactions via rule-based methods [15][16], while others capture statistical dependency between text elements in a sequence to predict the next element [17]. These earlier attempts rely heavily on either experts’ rules established prior or a set of probabilistic measures, scoping their capabilities within predefined scenarios. With the involvement of deep learning, the performance of generative models has been substantially improved, unlocking opportunities for a vast range of applications. Meanwhile, the critical need for trustworthiness in these models is also raised [18].
@@
The contributions of this paper are twofold. First, this paper offers an in-depth review of the historical development of AIGC, mapping its journey from early rule-based systems to advanced neural network methodologies. More importantly, this review addresses both technological advancements and challenges, such as scalability and ethical dilemmas.
Second, unlike most of existing review papers that focus on either a specific application domain, such as text-to-3D generation[19][20], or a particular type of methodology such as Video GANs[21][22], our paper resembles [23] in its exploration of the historical evolution of AIGC but with even broader coverage. More importantly, this paper integrates a wider range of methodologies and applications into a cohesive framework that spans multiple phases of AIGC development. A unique contribution of this work is the use of a consistent example throughout, which not only enables readers to follow the evolution of AIGC methods but also makes it easier to understand and compare the differences among various approaches within a practical context. While we share the same view of the challenges facing AIGC, such as bias, ethical concerns, and scalability of models [24], this paper extends the discussions by proposing research directions to tackle these issues. With these directions outlined, this paper contributes to the ongoing discourse on responsible AIGC development in real-world applications.
The rest of the paper is organized as follows. Section II provides the criteria for literature inclusion in this review. A detailed exploration of AIGC is given in Section III, presenting its core characteristics and advantages such as automation, consistency, and efficiency. Section IV delves into the early rule-based systems that lay the foundation for AIGC, followed by a discussion of the evolution towards statistical methods, deep learning models, and the integration of transfer and pre-trained models in Sections V, VI, and VII, respectively. Each section examines technological advancements and applications of these methodologies in real-world scenarios. Section VIII addresses the potential challenges and limitations of current AIGC technologies. Finally, Section IX concludes the review by summarizing the comprehensive insights gathered throughout the paper and suggesting future research directions.
@@ II Criteria for Inclusion in the Literature Review
To conduct a structured review of AIGC research, a clear literature search strategy is developed in alignment with our research objectives. The search begins with the identification of main keywords, including “Artificial Intelligence Generated Content,” “Generative AI,” and “Machine-generated Content”. Additional terms such as “transformer-based content generation” and “neural network content creation” are also incorporated to expand the scope of the relevant materials.
Six major academic databases are then selected for the search: IEEE Xplore, ACM Digital Library, ScienceDirect, SpringerLink, arXiv, and Google Scholar. To ensure comprehensive coverage of the AIGC’s historical development, additional sources are also considered, including MDPI, Frontiers, Wiley, Taylor and Francis.
The initial query returns thousands of articles. To manage this volume, a two-step screening process is applied:
  1. 1.
    Preliminary Screening. Two researchers review the title and abstract of each article independently. Papers are excluded if they are not peer-reviewed, not written in English, duplicated, or not directly relevant to AIGC—e.g., works that only mention generative tools without detailed elaboration of underlying models, algorithms, or core applications.
  2. 2.
    Full-Text Evaluation. After the initial screening, a total of 274 papers remain and then undergo full-text evaluation. Backward and forward citation searches are also performed to reduce the risk of omitting important studies. As part of the step, all papers are categorized into four developmental stages of AIGC evolution as mentioned in the Introduction: early rule-based systems, statistical models, deep learning models, and transfer learning and pre-trained models. Within each category, papers are reviewed chronologically to capture key milestones in the field’s progression and then selected based on the following two criteria:
    • Focused scope: The work explicitly develops or analyzes AIGC models or algorithms;
    • Technical depth: The work presents a clear and detailed explanation of model design, training procedures, and evaluation experiments.
For each milestone topic, if multiple articles are identified on the same subject, priority is always given to papers with higher citation counts or those published in journals with higher impact factors. When these papers are from the same research group, only the most foundational or representative work is retained to avoid redundancy. However, if they originate from different research groups, multiple works might be included, provided that each features a distinct technical perspective. The full screening process and the number of papers retained at each stage are presented in Fig. 2.
Based on the aforementioned literature search strategies, our review spans the past century, primarily encompassing all relevant publications released prior to 2024. Although the formal search window ends in 2023, a few key articles released in early 2024 are also included—particularly those that discuss emerging challenges and future directions in AIGC development. In total, 155 high-quality papers are selected for detailed analysis in this review. For each of them, key information is extracted, including the generative task type, model architecture, datasets, reported limitations, and suggestions for future work.
To validate our search process, VOSviewer is utilized to visualize keyword co-occurrence analysis, as shown in Fig. 3. It is clear that the key terms are well-clustered into several coherent groups, highlighting the comprehensiveness and relevance of our retrieved literature. For instance, the red cluster focuses on the foundational models and methodological underpinnings of AIGC, with keywords like ”deep neural network,” ”Transformer,” ”BERT,” ”deep learning,” ” transfer learning,” and ”fine-tuning.” The blue cluster emphasizes generative approaches, such as ”autoencoder,” ”latent space,” ”synthesis”, and ”diffusion model.” Meanwhile, the green cluster centers around general AI and content generation applications, including keywords like ”algorithm,” ”feature”, ”image,” ”video”, and ”audio”. Overall, the structured distribution of terms with the presence of both foundational and applied keywords confirm that our search strategies successfully captured a diverse and balanced spectrum of relevant literature.
Finally, the distribution of all selected articles by publication year is presented in Fig. 4, illustrating the temporal trends in the development of AIGC-related research. As illustrated, the number of publications on AIGC has increased significantly in recent years, particularly after 2020. This trend coincides with the emergence and widespread adoption of transformer-based models and large-scale generative frameworks (e.g., GPT-4[1]). The sharp rise in 2022 and 2023 suggests a growing academic interest and a transition from foundational research to application-oriented studies across multiple modalities, including text, image, audio, and video generation.
Figure 2: Literature search and screening process
Figure 3: Keyword co-occurrence networks after screening.
@@ III General AIGC Technique Framework
AIGC is an innovative way of autonomously generating diverse forms of content by algorithms, machine learning models, or other artificial intelligence techniques. The main characteristic of AIGC is its reliance on data input and pre-generated models to produce content, requiring less or no human intervention. Compared to manual creation, AIGC has distinguished advantages in terms of automation, scale, and efficiency [25]. It also ensures consistency, customization capability [26], and cost-effectiveness [27]. These strengths have continuously motivated researchers [28] to advance AIGC technologies for decades, leading to a variety of applications as shown in Table I. Beginning with basic text generation [29], AIGC has now expanded to include video, audio [30], and 3D model [31] generation for applications in fields such as education [32], healthcare [33], entertainment [34], and many others.
@@
The flow of an AIGC system begins with user input that goes through a pre-processing phase to convert from its raw, naturally human-understandable form into a structured, machine-interpretable format [66]. This usually involves techniques such as natural language processing (NLP) [67] for text data or feature extraction for visual [68] or auditory data, with methods like decision trees and pattern matching bridging input rules to algorithms.
@@
These sequential steps, each serving a unique role, contribute collaboratively to the robustness and efficacy of the content generation process. A deep understanding of these steps in each developmental milestone of AIGC helps us gain insights into where AIGC has been and where it is heading. To that end, this paper uses the same example of “generate a research question with the keywords: artificial intelligence, healthcare, and ethical implications” across all milestones. This unified reference point not only demonstrates the unique advantages of the methodologies at each milestone, such as the transparency of rule-based systems, the data-driven nature of statistical methods, the automation capabilities of deep learning, and the flexibility of transfer learning, but also makes comparisons between approaches more straightforward and clear. Ultimately, the paper aims to assist researchers in better grasping each method’s operational process, enabling them to choose a proper one for their practical applications or combine different ones for optimal results.
The choice of text generation as the recurring example stems from natural language generation’s earlier development and structured theoretical lineage. Considering the conceptual overlaps among generative models and frameworks across various modalities, especially in how they manage data representations and content synthesis, highlighting text generation serves to leverage a mature example for clarity, without narrowing the scope of our review to text alone.
@@ IV Early Rule-based Systems
The roots of AIGC can be traced back to the 1950s [71][72] when early rule-based systems lay the foundation for its development. These systems employ a set of expert rules defined a priori to guide content generation. Figure 6 presents the specific framework for rule-based systems within the context of AIGC discussed in the previous section.
@@ IV-A Fundamentals
Early rule-based systems primarily focus on texts, using NLP techniques to enable computers to understand human inputs, in which the preprocessing step is attributed to ‘parsing’. The common methods include tokenization, stemming, stop word removal, rule-based parsing, and entity recognition. Tokenization breaks down the input data into small pieces, like words. Doing so converts a continuous stream of text into a manageable sequence of tokens, making it easier for algorithms to interpret. Stemming reduces words to their basic form, like ’running’ to ’run’. This ensures consistency across similar words, reducing complexity with fewer unique words. In the step of stop word removal, common yet nonessential words like ’the’ and ’is’ are removed. This step helps the system focus on meaningful words only and reduces noise in the data. Rule-based parsing applies grammar rules to understand sentence structure, aiding in the interpretation of word relationships and the comprehension of text syntax and semantics. In addition, Entity recognition helps spot and categorize keywords, enabling the quick identification of important information such as the text’s main subjects. These steps streamline the complexity of natural language, making it more structured and manageable for analysis and encoding.
@@
The role of post-processing in the rule-based phase is pretty straightforward. It usually consists of some content enrichment processes, such as sentence re-organization, spelling, and grammar corrections, to improve the clarity and fluency of generated results. The decision to enhance the consistency or diversity, and depth of the language used is made through content formatting and lexical enrichment.
@@ IV-B Example
Early rule-based systems find applications in text generation. For example, Eliza is designed as a chatbot for psychotherapist conversations, SHRDLU advances semantic analysis, and Tale-Spin [76] can create poetry and stories. Here, we detail how Eliza responds to the user input “generate a research question with the keywords: artificial intelligence, healthcare, and ethical implications” to showcase the operations of early rule-based systems.
@@ V-A Fundamentals
Gaussian Mixture Models (GMM) 1980 L.P. Hansen[79]
@@
In the realm of statistical language processing, N-gram models serve as a straightforward yet effective method for predicting the next item in a sequence based on the preceding N items. They are particularly valuable for tasks such as lexical selection and modifier ordering [82] due to their ability to capture local context. Yet N-gram is effective for text generation and speech recognition [83], it often faces several limitations, especially when dealing with long-range dependencies and language ambiguity, owing to its fixed context window. To address these limitations, researchers develop Hidden Markov Models (HMMs), which build upon N-grams by incorporating hidden states, allowing for more flexible sequence modeling. HMMs are crucial for handling temporal variations, making them indispensable in tasks like part-of-speech tagging and speech recognition [84][85]. Nonetheless, HMMs are still constrained by their assumptions about state independence and often require extensive training data to achieve high accuracy.
@@ VI Deep Learning Methods
Neural networks and deep learning models mark a major leap forward in AIGC [90], overcoming the limitations of statistical methods by better managing the growing volume, dimensionality, and complexity of modern data, especially unstructured data such as images, text, and speech. Mimicking the structure of the human brain, neural networks consist of layers of interconnected nodes (neurons). As data flows through the layers, the networks conduct a step-by-step inference process, automatically recognizing patterns and making predictions from input to output. Extended from this foundational structure, deep learning uses convolutional neural networks (CNNs), in which the nodes at each layer are organized into overlapping clusters so that each cluster feeds data to multiple nodes in subsequent layers. This design enhances the network’s capability to process high-dimensional data to learn and capture complex relationships more effectively. As shown in Figure 11, deep-learning-based AIGC shares similarities with statistical methods in terms of feature engineering for user input. What sets them apart is the process of selecting, modifying, and creating features. Deep learning automates that process through training on diverse and large datasets, while statistical methods require human intervention. Notably, the quality and diversity of datasets are crucial for training deep learning models, as the data needs to represent a wide range of scenarios within the problem’s scope. Thus, data augmentation is often applied to enhance both the size and variability of the training dataset, improving the model’s generalization ability. In deep learning, the backbone architectures include Recurrent Neural Networks (RNNs), CNNs, and Transformer [91] as detailed in the following subsections.
@@ VI-C Transformer
As shown in Figure 12, the transformer consists of two main parts: multiple layers $j$ (i.e., $j=N$) of an encoder and a decoder that have very similar structures but serve different purposes. The former is responsible for encoding the input sequence into a context-rich representation, while the latter is used to generate new information in a sequence based on the given input and its encoded representation.
Before the input sequence is fed into the encoder, it is first tokenized, similar to that in early rule-based and statistical methods. What sets the transformer apart is the embedding process, mapping each token into a dense vector representation with a dimension of $d_{\text{model}}$. Note that such dense vectors are generated from an embedding matrix, either derived from a pre-trained model (e.g., BERT [94], GPT) or trained from scratch during the model’s training stage. As stated earlier, the transformer model processes the entire input in parallel, and it does not have an inherent understanding of the sequence order of the tokens. Therefore, positional encoding is essential, as it is ultimately added to the input embedding to help the model gain the necessary sequence order information. While sine and cosine are the most traditional and widely used forms of positional encoding, many other approaches are available, such as learned embeddings and rotary positional embeddings[95], and their choice highly depends on the model architecture and the task requirements.
@@
While initially introduced for NLP tasks, the modularity, and extensibility of the transformer model have made it a popular and powerful tool across a wide range of applications, such as modern large language models. Beyond excelling in text generation, it has recently been successfully extended to the area of computer vision, such as Vision Transformer [96], showcasing its effectiveness in vision generation.